Please help me solve this. I have the answer, but I'm stuck on how to work the problem to make my answer match the one that was given to me. I was told to use it as a guide to compare my work to the answer that was given to me and I'm really stuck. The problem is:
1/3x+2/1y=3/1
2/3x-3/5y=2/5

Answer by stanbon(75887)

(1/3)x + (2/1)y = (3/1)
(2/3)x - (3/5)y = (2/5)
Multiply the 1st equation by 2 to get:
(2/3)x + (4/1)y = (6/1)
Subtract the 2nd equation from that to solve for y, as follows:
(23/5)y = 28/5
y = 28/23
----------------
Substitute into one of the original equations to solve for x, as follows:
(1/3)x + 56/23= 3
(1/3)x = 69/23-56/23
(1/3)x = 13/23
x = 39/23
================
Cheers,
Stan H.